# Literature Review: Cyber Security Monitoring in Maritime

### v1

@@ III-B3 Data Extraction and Analysis
Following the SLR protocol, at this stage, the papers were studied deeply, summarised, and assessed. The data extracted from each paper were:
@@
  • Introduced monitoring method.
  • Data type processed by the monitoring solution.
  • Data and implementation availability.
  • Performance and computational cost analysis.
  • Validation/evaluation of the presented results.
@@ IV Bibliometric analysis
In light of the information presented, the temporal dissemination of the identified papers is introduced in Figure 1. The annual distribution of published papers does not reveal a discernible trajectory. However, an interesting observation comes up: the number of publications started rising after the major cyber incidents in 2017 within the maritime sector (e.g., the cyber attack against Maersk described in Section I).
Figure 1: Year of publication.
@@
As can be seen from Figure 1, 21 research papers out of the 23 originate from 2018–2023, whereas only two research papers have been published before 2018. In other words, maritime security monitoring is a relatively new research domain which has started to receive more attention during the last 5–6 years.
@@
In accordance with the representation delineated in Table II, the preponderance of contributions assumes the form of conference papers, while about a third of the corpus manifests as journal articles. As for conference papers, it is worthwhile to note that the vast majority of the papers (14 in total) have been published by IEEE, whereas only one paper originates from an ACM conference. When investigating the conference venues more closely, we found that three papers were published at the International Conference on Cyber Situational Awareness, Data Analytics and Assessment, whereas each remaining conference accounted for one paper. As for journal papers, we found that three papers appeared in IEEE Transactions on Intelligent Transportation Systems. Furthermore, Ocean Engineering, Applied Ocean Research, MDPI Journal of Maritime Science and Engineering, MDPI Information, and MDPI Energies published one paper each.
TABLE II: Publications’ type and publisher
Type Total IEEE ACM Elsevier MDPI
Journal article 8 3 - 2 3
Conference paper 15 14 1 - -
Total 23
The 23 research papers were authored by 99 researchers, and we also tried to identify more prolific authors in the field. For that task, we used a frequent itemset mining algorithm to identify all authors who have published at least two papers. Since the purpose of the frequent itemset mining algorithm is to identify associations, that allowed us to detect not just single authors, but also prolific co-author groups. The results are shown in Table III.
TABLE III: Prolific authors
Authors Number of papers
Boudehenn, Boudraa, Cexus 2
Boudvin, Brosset, Kermarrec 2
Jacq 2
Laso 2
@@
As Table III indicates, no author has published more than two papers in the research domain, and a small minority of authors (8 out of 99) have published more than one paper. That indicates that for the majority of authors, maritime security monitoring has not been a persistent focus area with a larger number of peer-reviewed academic publications. In a recent SLR paper by Bolbot et al. [3], a similar phenomenon has been reported – only 17% of authors have published more than one paper in the field of maritime cyber security.
Also, although we identified two co-author groups in Table III, we did not find any group with a substantial number of published papers. Furthermore, as two co-author groups from Table III are coming from the same country, we did not detect any indications of prolific international research cooperation. This illustrates the fact that maritime security monitoring has received modest attention so far, and well-established domestic and international research groups specifically focusing on this particular field have not yet emerged.
@@ V-A Taxonomy
To answer RQ2 (see Section III-A), we have created a taxonomy for existing research literature that is presented in Figure 2. Table IV provides detailed information about the selected 23 research papers and how they map to different taxonomy categories. As Figure 2 illustrates, the research papers can be categorised by the environment they target, the type of data which the proposed methods analyse, the type of the proposed methods, and the type of cyber attacks and anomalies that the proposed methods are able to detect. Note that these four categorisations from Figure 2 correspond to the “Targeted Environment”, “Monitored Data”, “Method Type”, and “Detected Attacks and Anomalies” columns in Table IV.
Figure 2: Taxonomy for existing research literature.
TABLE IV: Overview of Research Papers
@@
Type Environment Type Attacks
and
@@
[36] 2023 M navigational system application layer protocol unsupervised ensemble of spoofing
(NMEA GPS) unsupervised methods
@@
[15] 2021 M navigational system application layer protocol unsupervised One-Class Support spoofing
(NMEA GPS) Vector Machine
@@
[44] 2020 M vessel CPS OT (ship propulsion system) graph-based graph-based method CPS attacks
@@
[49] 2018 M navigational system application layer protocol rule-based distributed rule- spoofing
(AIS) based processing
@@
SCADA network) without any specific
monitoring method
@@
When using Table IV for analysing the mapping of papers to taxonomy categories in Figure 2, we discovered four non-overlapping clusters of closely related papers. Each cluster contains five studies and the clusters are thus covering 20 papers out of 23. The clusters have been highlighted with different colours (red, blue, green, and orange) in Figure 2. If the majority of papers in some category belong to a single cluster, the category is highlighted with the colour of the respective cluster in Figure 2. Note that the “Distributed maritime system or network on land and sea” category in Figure 2 is highlighted with two colours, since it contains papers from two clusters which cover 50% and 30% of that category, respectively. Also, the “Unsupervised” category in Figure 2 is not highlighted with any colour, since it is the only category which contains papers from three clusters, and each cluster covers less than half of the papers in that category. The four clusters from Figure 2 are summarised below:
  • Studies on security monitoring for navigational systems [36, 40, 15, 45, 49] (red cluster in Figure 2) – papers which propose the analysis of application layer protocol data for detecting AIS and GPS spoofing attacks and anomalies in navigational systems.
@@
  • Studies on security monitoring for distributed maritime systems and networks [34, 35, 37, 38, 53] (green cluster in Figure 2) – papers which propose the analysis of network traffic for detecting attacks against maritime IoT networks, maritime Software Defined Networks (SDNs), and navy networks.
@@
  • Studies on security monitoring architectures and testbeds [14, 39, 47, 13, 48] (orange cluster in Figure 2) – the papers which have not proposed one particular security monitoring method like the studies from previous clusters, but rather a security monitoring architecture or testbed.
@@
Detected clusters from Figure 2 reveal some associations between taxonomy categories which describe the nature of existing research. First, studies on security monitoring for navigational systems (red cluster) have focused on the analysis of relevant application layer protocols, and rule-based security monitoring approaches have been usually applied in the context of navigational systems. Furthermore, studies on security monitoring for distributed maritime systems and networks (green cluster) have focused on network traffic analysis, and such studies tend to employ supervised ML methods. Also, studies on security monitoring for CPSs (blue cluster) have proposed the analysis of OT data, and graph-based approaches have been employed only in this particular context. Unlike papers from other clusters, the studies on security monitoring architectures and testbeds (orange cluster) propose a range of different methods and their simultaneous use with the ability to process different types of input data and detect different attacks and anomalies.
@@
It should be pointed out that three papers not covered by the above clusters are fairly close to them. Although in [50, 52] the authors have described an architecture and a testbed for processing OT data only, these papers are otherwise similar to the orange cluster in Figure 2. Also, although the paper [42] focuses on the analysis application layer data instead of network traffic, it proposes a security monitoring method for maritime IoT systems, and is thus closely linked to the green cluster in Figure 2. In the remainder of the paper, we will consider these three papers together with their closest clusters.
@@
In order to distinguish papers that have proposed security monitoring methods from the papers that have described architectures and testbeds, we divided all papers by their type (see the “Paper Type” column in Table IV). Letter M denotes the papers that discuss specific security monitoring methods, letter A the papers which describe the architecture of the entire security monitoring system, and letter T the papers that discuss testbeds designed for security monitoring data collection, testing, and exercises. As can be seen from Table IV, 16 papers have focused on specific security monitoring methods [43, 34, 35, 49, 53, 15, 41, 44, 36, 37, 42, 40, 45, 51, 38], while 4 papers have described architectures [13, 48, 47, 50] and 3 papers have described testbeds [52, 14, 39], respectively. Most papers that concern architectures and testbeds [13, 52, 48, 47, 14, 39] have not proposed any novel security monitoring algorithms for analysing specific data (e.g., network traffic), but have rather employed various well-known industrial security monitoring technologies such as signature-based network IDS, antivirus, etc. which monitor a wide range of different data types (see the “Monitored Data” and “Monitoring Method” columns in Table IV). Therefore, we have set the “Method Type” to architecture or testbed for these papers in Table IV, because the focus of the paper lies on the entire architecture or testbed, not on one security monitoring method. The only exception is the work by Jia et al. [50] which describes the use of an Isolation Forest-based algorithm within the proposed architecture.
@@
As can be seen from Table IV, security monitoring architectures received a significant amount of attention until 2019 (all 4 relevant papers were published in 2018–2019, constituting almost half of the 9 papers published before 2020). In contrast, during the more recent years after 2019, the research focus has shifted towards security monitoring methods and testbeds, with the majority of the papers (12 out of 14) discussing security monitoring methods. That finding illustrates the fact that during the earlier research period before 2020, scholars preferred to focus on the architecture of the entire security monitoring system, studying the unique challenges that arise in maritime environments. After the architectural issues and solutions for them were identified, the attention moved to particular security monitoring methods which constitute the building blocks of security monitoring system architectures.
@@
The “Targeted Environment” column in Table IV indicates that most papers have proposed methods, architectures, and testbeds for addressing the needs of vessels or a distributed environment which involves vessels and land-based objects. In contrast, a small fraction of works target land-based environments only [46, 47].
@@
As for the nature of monitored data (see the “Monitored Data” column in Table IV), we have used the term network traffic for studies which have proposed network monitoring methods without considering the application layer protocol in network packets. In the rest of the article, we will call such methods network traffic monitoring methods. Almost all the studies published until 2022 were not focusing on monitoring network traffic, with the work by Ptasinski et al. [53] being the only exception. For 2023, a new trend can be observed where a significant proportion of papers have described such methods [34, 35, 37, 38]. However, these studies have been limited to specific environments like maritime IoT networks and SDN systems. Also, according to Figure 2, the majority of papers that have proposed the analysis of application layer data have employed it for security monitoring of navigational systems, whereas the analysis of OT data is generally used in the context of CPS.
@@
As for the types of security monitoring methods (see the “Method Type” column in Table IV), supervised and unsupervised ML approaches are the most common, while some papers have also described rule-based and graph-based approaches. As for supervised ML, deep learning-based approaches have been the most widely employed [34, 46, 37, 38], whereas no unsupervised ML method has seen more frequent use than the others. As Figure 2 illustrates, more than half of the papers which have proposed supervised ML approaches have applied them for maritime network security monitoring, whereas rule-based approaches tend to be applied for navigational systems. The only paper which has described a graph-based approach [44] has employed it for CPS security monitoring.
TABLE V: Overview of Research Papers on Security Monitoring Methods
@@
Paper Evaluation Is Data Is Data Is Code Performance Computational Detected
@@
[34] NSL-KDD no yes no yes no network attacks against maritime IoT systems
[35] CICDDoS2019 no yes no yes yes DDoS attacks against maritime IoT systems
[36] MARSIM yes yes yes yes no GPS spoofing
[37] TON-IoT no yes no yes no network attacks against maritime IoT systems
[38] X-IIoTID no yes no yes no network attacks against
@@
[40] internal yes no no no no anomalies in NMEA messages
[15] internal yes no no yes yes GPS spoofing
[41] internal yes no no no no attacks against ship’s CPSs
[42] DS2OS no yes no yes no application layer attacks against
@@
[43] internal yes no no no no attacks against ship’s CPSs
[44] internal yes no no yes no attacks against ship’s CPSs
[45] internal yes no no no no anomalies in AIS messages
[46] internal yes no no no no attacks against canal lock CPS
[49] MarineTraffic AIS yes no no yes no AIS spoofing
[50] internal yes no no no no anomalies in the state and trajectory
@@
[51] internal yes no no no no anomalies in ship’s CPSs
[53] internal yes no no no no network flooding attacks against navy networks
@@
Table V provides an overview of 17 papers that propose security monitoring methods (i.e., the papers from Table IV with the “Method Type” column containing other values than architecture or testbed). Similarly to Table IV, Table V contains the “Detected Attacks and Anomalies” column, but provides more detailed information about the nature of attacks and anomalies. As this column indicates, some studies have proposed methods for detecting specific attack types. For example, whereas methods from [34, 37, 38] can detect network attacks of various types, algorithms from [35, 53] detect Distributed Denial of Service (DDoS) and network flooding attacks only. As another example, [40] proposes generic anomaly detection for NMEA messages, while [15, 36] focus on processing NMEA messages for the detection of GPS spoofing. Other methods described in papers from Table V include AIS anomaly detection [45], detection of AIS spoofing [49], detection of application layer attacks against maritime IoT systems [42], anomaly detection in the state and trajectory of unmanned surface vehicles [50], detection of attacks against canal lock CPS [46], and detection of attacks and anomalies in the CPSs of ships [43, 41, 44, 51].
@@
Table V also provides some insights into the evaluation process of the proposed security monitoring methods. As can be seen from Table V, whenever maritime-specific data has been used for evaluating the method, the dataset is generally private (i.e., not publicly available). Also, whenever a public dataset has been used for evaluation, it is generally not maritime-specific. The only exception is the study by Spravil et al. [36], where the authors have created and released the MARSIM dataset.
@@
Table V also reveals several other aspects of method evaluations – not all studies have published widely acknowledged performance metrics like precision, recall, and F1-score which provide a realistic picture of the method performance (for such studies, the “Performance Evaluation” column in Table V contains no). Also, most studies have not released the implementation of the method (see the “Is Code Publicly Available” column in Table V) and have not assessed the computational cost of the method (see the “Computational Cost Assessment” column in Table V).
@@
The following sections provide a more detailed discussion of all papers from Table IV. For structuring the discussion, we have closely followed the aforementioned clustering of the selected research papers into four clusters based on the taxonomy presented in Figure 2. The papers on security monitoring architectures and testbeds (with the “Paper Type” column set to A or T in Table IV) are covered in Sections V-E and V-F, respectively (note that such papers correspond to the orange cluster in Figure 2). The discussion of the remaining papers on security monitoring methods (with the “Paper Type” column set to M in Table IV) has been divided into three subsections by the targeted environment (represented by red, blue, and green clusters in Figure 2) – the papers concerning security monitoring methods for distributed maritime systems and networks are covered in Section V-B, papers on security monitoring methods for navigational systems are covered in Section V-C, and papers on security monitoring methods for maritime CPSs are covered in Section V-D.

### v2

@@ III-B3 Data Extraction and Analysis
Following the SLR protocol, at this stage, the papers were fully assessed and relevant data were extracted relating to:
@@
  • The nature of the proposed security monitoring method.
  • The nature of the maritime environment the proposed method was designed for.
  • The nature of the data processed by the proposed method.
  • The nature of the cyber attacks and anomalies detected by the proposed method.
  • Performance evaluation details for the proposed method.
  • The nature of the performance evaluation dataset and its availability information.
  • Computational cost analysis for the proposed method.
  • Availability information for the implementation of the proposed method.
@@ IV Bibliometric analysis
In Figure 1, the annual distribution of the identified papers is shown, which does not reveal a clear pattern. However, an interesting observation can be made: the number of publications started rising after the major cyber incidents in 2017 within the maritime sector (e.g., the cyber attack against Maersk described in Section I). As can be seen from Figure 1, 30 research papers out of the 32 originate from 2018–2025, whereas only two research papers have been published before 2018. In other words, maritime security monitoring is a relatively new research domain which has started to receive more attention during the last 6–7 years.
Figure 1: Publications per year
@@
According to Table II, conference papers form a slight majority of the literature. On the other hand, journal articles are dominant among recent papers from 2023–2025. This indicates the increasing quality of relevant research. The most prominent publisher is IEEE with 24 papers (i.e., 75% of the analysed literature). As for conference papers, the vast majority of them (17 out of 18) have been published by IEEE, whereas only one paper originates from an ACM conference. When investigating the conference venues more closely, we found that three papers were published at the International Conference on Cyber Situational Awareness, Data Analytics and Assessment, whereas each remaining conference accounted for one paper. As for journal papers, three papers appeared in IEEE Transactions on Intelligent Transportation Systems and two papers in IEEE Access. Remaining journals published one paper each.
TABLE II: Publications per publisher
Publisher Journal Conference
article paper
IEEE 7 17
ACM - 1
Elsevier 3 -
MDPI 3 -
Springer 1 -
Total 14 18
@@
The 32 research papers were authored by 126 researchers, and we also tried to identify more prolific authors in the field. For that task, we used a frequent itemset mining algorithm to identify all authors who have published at least two papers. Since the purpose of the frequent itemset mining algorithm is to identify associations, that allowed us to detect not just single authors, but also prolific co-author groups. According to our findings, no author has published more than three papers in this research domain, and a small minority of authors (16 out of 126 or 12.7%) have published more than one paper. That indicates that for the majority of authors, maritime security monitoring has not been a persistent focus area with a larger number of peer-reviewed academic publications. In a recent SLR paper by Bolbot et al. [3], a similar phenomenon has been reported – only 17% of authors have published more than one paper in the field of maritime cyber security. Also, we identified one co-author group with three publications, with the co-authors coming from the same country. This illustrates the fact that maritime security monitoring has received limited attention so far, and well-established prolific research groups specifically focusing on this particular field have not yet emerged.
@@ V-A Taxonomy
To answer RQ2 (see Section III-A), we have created a taxonomy for existing research literature that is presented in Figure 2. Table III provides detailed information about the selected 32 research papers and how they map to different taxonomy categories. As Figure 2 illustrates, the research papers can be categorised by the environment they target, the type of data the proposed methods analyse, the type of the proposed methods, and the type of cyber attacks and anomalies that the proposed methods are able to detect. Note that these four categorisations from Figure 2 correspond to the “Targeted Environment”, “Monitored Data”, “Method Type”, and “Detected Attacks and Anomalies” columns in Table III.
Figure 2: Taxonomy for existing research literature
TABLE III: Overview of research papers
@@
Type Environment Type Attacks and
@@
[30] 2025 M navigational system application layer protocol supervised deep learning spoofing and
(Navico BR24 and NMEA) data manipulation
[31] 2025 M distributed maritime network traffic supervised deep learning IoT attacks
system/network
[32] 2025 M distributed maritime application layer protocol other stochastic IoT attacks
system/network (NMEA) semi-Markov processes
[33] 2024 M navigational system application layer protocol other semi-supervised spoofing and
(Navico BR24 and NMEA) learning data manipulation
[34] 2024 M distributed maritime network traffic supervised CatBoost network
system/network layer attacks
[35] 2024 M distributed maritime network traffic supervised, deep learning, network
system/network unsupervised Isolation Forest layer attacks
[36] 2024 T navigational system application layer protocol testbed various (different spoofing and
(Navico BR24 and NMEA) types of IDS) data manipulation
[37] 2023 M navigational system application layer protocol rule-based policy enforcement spoofing and
(ASTERIX) data manipulation
[38] 2023 T distributed maritime various (network traffic, testbed various (IDS signatures, various
system/network NMEA, etc.) SIEM queries, etc.)
@@
[41] 2023 M navigational system application layer protocol unsupervised ensemble of spoofing and
(NMEA GPS) unsupervised methods data manipulation
@@
[15] 2021 M navigational system application layer protocol unsupervised One-Class Support spoofing and
(NMEA GPS) Vector Machine data manipulation
@@
[49] 2020 M vessel CPS OT (ship propulsion system) other graph-based method CPS attacks
@@
[54] 2018 M navigational system application layer protocol rule-based distributed rule- spoofing and
(AIS) based processing data manipulation
@@
SCADA network) without any specific method
@@
When using Table III for analysing the mapping of papers to taxonomy categories in Figure 2, we discovered four non-overlapping clusters of closely related papers. The clusters cover 30 papers out of 32, and have been highlighted with different colours (red, blue, green, and orange) in Figure 2. If the majority of papers in some category belong to a single cluster, the category is highlighted with the colour of the respective cluster in Figure 2. Note that the “Unsupervised” and “Other methods” categories in Figure 2 are not highlighted with any colour, since there is no cluster which would cover the majority of papers in these categories. The four clusters from Figure 2 are summarised below:
  • Studies on security monitoring for navigational systems [41, 45, 15, 50, 54, 33, 37, 30] (red cluster in Figure 2) – papers which propose the analysis of application layer protocol data for detecting AIS and GPS spoofing attacks, radar image manipulation attacks, and anomalies in navigational systems.
@@
  • Studies on security monitoring for distributed maritime systems and networks [39, 40, 42, 43, 58, 35, 31, 34] (green cluster in Figure 2) – papers which propose the analysis of network traffic for detecting attacks against maritime IoT networks, maritime Software Defined Networks (SDNs), maritime wireless networks, and navy networks.
@@
  • Studies on security monitoring architectures and testbeds [14, 44, 52, 13, 53, 57, 55, 36, 38] (orange cluster in Figure 2) – papers which have not proposed one particular security monitoring method like the studies from previous clusters, but rather a security monitoring architecture or testbed.
@@
The identified clusters in Figure 2 reveal some associations between taxonomy categories which describe the nature of existing research. First, studies on security monitoring for navigational systems (red cluster) have focused on the analysis of relevant application layer protocols, and rule-based security monitoring approaches have been usually applied in the context of navigational systems. Furthermore, studies on security monitoring for distributed maritime systems and networks (green cluster) have focused on network traffic analysis, and such studies tend to employ supervised ML methods. Also, studies on security monitoring for CPSs (blue cluster) have proposed the analysis of OT data, and there is no particular method category strongly associated with this cluster. Unlike papers from other clusters, the studies on security monitoring architectures and testbeds (orange cluster) propose a range of different methods and their simultaneous use with the ability to process different types of input data and detect different attacks and anomalies.
@@
It should be pointed out that two papers, [47, 32], which are not covered by the above clusters are fairly close to one of them. Although these two papers focus on the analysis of application layer data instead of network traffic, they propose security monitoring methods for maritime IoT systems, and are thus closely linked to the green cluster in Figure 2. In the remainder of our study, we will consider these two papers together with their closest cluster.
@@
In order to distinguish papers that have proposed security monitoring methods from the papers that have described architectures and testbeds, we divided all papers by their type (see the “Paper Type” column in Table III). Letter M denotes the papers that discuss specific security monitoring methods, letter A the papers which describe the architecture of the entire security monitoring system, and letter T the papers that discuss testbeds designed for security monitoring data collection, testing, and exercises. As can be seen from Table III, 23 papers have focused on specific security monitoring methods [48, 39, 51, 40, 54, 58, 15, 46, 49, 41, 42, 47, 45, 50, 56, 43, 35, 33, 32, 31, 37, 34, 30], while 4 papers have described architectures [13, 53, 52, 55] and 5 papers have described testbeds [57, 14, 44, 36, 38], respectively. Most papers that concern architectures and testbeds [13, 57, 53, 52, 14, 44, 36, 38] have not proposed any novel security monitoring algorithms for analysing specific data (e.g., network traffic), but have rather employed various well-known industrial security monitoring technologies such as signature-based network IDS, antivirus, etc., which monitor a wide range of different data types (see the “Monitored Data” and “Monitoring Method” columns in Table III). Therefore, we have set the “Method Type” to architecture or testbed for these papers in Table III, because the focus of these papers lies on the entire architecture or testbed, not on one security monitoring method. The only exception is the work by Jia et al. [55] which describes the use of an Isolation Forest-based algorithm within the proposed architecture.
@@
As can be seen from Table III, security monitoring architectures received a significant amount of attention until 2019 (all 4 relevant papers were published in 2018–2019, constituting almost half of the 9 papers published before 2020). In contrast, during the more recent years after 2019, the research focus has shifted towards security monitoring methods and testbeds, with the majority of the papers (19 out of 23) discussing security monitoring methods. This finding illustrates the fact that during the earlier research period before 2020, scholars preferred to focus on the architecture of the entire security monitoring system, studying the unique challenges that arise in maritime environments. After the architectural issues and solutions for them were identified, the attention moved to particular security monitoring methods which constitute the building blocks of security monitoring system architectures.
@@
The “Targeted Environment” column in Table III indicates that most papers have proposed methods, architectures, and testbeds for addressing the needs of vessels or a distributed environment which involves vessels and land-based objects. In contrast, a small fraction of works target land-based environments only [51, 52].
@@
As for the nature of monitored data (see the “Monitored Data” column in Table III), we have used the term network traffic for studies which have proposed network monitoring methods without considering the application layer protocol in network packets. In the rest of the article, we will call such methods network traffic monitoring methods. Almost all the studies published until 2022 were not focusing on monitoring network traffic, with the work by Ptasinski et al. [58] being the only exception. Starting from 2023, a new trend can be observed where a significant proportion of papers have described such methods [39, 40, 42, 43, 35, 31, 34]. However, these studies have been limited to specific environments like maritime IoT networks. Also, according to Figure 2, the majority of papers that have proposed the analysis of application layer data have employed it for security monitoring of navigational systems, whereas the analysis of OT data is generally used in the context of CPS.
@@
As for the types of security monitoring methods (see the “Method Type” column in Table III), supervised and unsupervised ML approaches are the most common, while some papers have also described rule-based and other approaches. As for supervised ML, deep learning (DL) based approaches have been the most widely employed [39, 51, 42, 43, 30, 31, 35], whereas no unsupervised ML method has seen more frequent use than the others. As Figure 2 illustrates, more than half of the papers which have proposed supervised ML approaches have applied them for maritime network security monitoring, whereas rule-based approaches tend to be applied for navigational systems.
TABLE IV: Overview of research papers on security monitoring methods
@@
Paper Year Evaluation Is Data Is Data Is Code Performance Computational Detected
@@
[30] 2025 RadarPWN yes yes no yes no radar image manipulation attacks
[31] 2025 SOLIDS yes yes no yes yes DDoS attacks against
maritime underwater IoT systems
[32] 2025 internal yes no no yes no DDoS attacks against
maritime IoT systems
[33] 2024 RadarPWN yes yes no yes no radar image manipulation attacks
[34] 2024 internal yes no no yes yes network attacks against
navy satellite networks
[35] 2024 WSN-DS no yes no yes no DoS attacks against
maritime wireless networks
[37] 2023 internal yes no no yes yes radar image manipulation attacks
[39] 2023 NSL-KDD no yes no yes no network attacks against
maritime IoT systems
[40] 2023 CICDDoS2019 no yes no yes yes DDoS attacks against
maritime IoT systems
[41] 2023 MARSIM yes yes yes yes no GPS spoofing
[42] 2023 TON-IoT no yes no yes no network attacks against
maritime IoT systems
[43] 2023 X-IIoTID no yes no yes no network attacks against
@@
[45] 2022 internal yes no no no no anomalies in NMEA messages
[15] 2021 internal yes no no yes yes GPS spoofing
[46] 2021 internal yes no no no no attacks against ship’s CPSs
[47] 2021 DS2OS no yes no yes no application layer attacks against
@@
[48] 2020 internal yes no no no no attacks against ship’s CPSs
[49] 2020 internal yes no no yes no attacks against ship’s CPSs
[50] 2020 internal yes no no no no anomalies in AIS messages
[51] 2019 internal yes no no no no attacks against canal lock CPS
[54] 2018 MarineTraffic yes no no yes no AIS spoofing
AIS
[55] 2018 internal yes no no no no anomalies in the state and trajectory
@@
[56] 2018 internal yes no no no no anomalies in ship’s CPSs
[58] 2013 internal yes no no no no network flooding attacks against
navy networks
@@
Note that unlike the papers on architectures and testbeds, the studies of security monitoring methods involve thorough evaluations of particular algorithms on dedicated datasets. In order to get more insights into the relevant literature, Table IV provides an overview of 24 papers that propose security monitoring methods (i.e., the papers from Table III with the “Method Type” column containing other values than architecture or testbed). Similarly to Table III, Table IV contains the “Detected Attacks and Anomalies” column, but provides more detailed information about the nature of attacks and anomalies. As this column indicates, some studies have proposed methods for detecting specific attack types. For example, whereas methods from [39, 42, 43] can detect network attacks of various types, algorithms from [40, 58, 32, 35, 31] are focusing on DDoS/DoS and network flooding attacks only. As another example, [45] proposes generic anomaly detection for NMEA messages, while [15, 41] focus on processing NMEA messages for the detection of GPS spoofing. Other methods described in papers from Table IV include detection of radar image manipulation attacks [33, 37, 30], AIS anomaly detection [50], detection of AIS spoofing [54], detection of application layer attacks against maritime IoT systems [47], anomaly detection in the state and trajectory of unmanned surface vehicles [55], detection of attacks against canal lock CPS [51], and detection of attacks and anomalies in the CPSs of ships [48, 46, 49, 56].
@@
Table IV also provides some insights into the evaluation process of the proposed security monitoring methods. In older studies which used maritime-specific data for evaluations, datasets were generally private (i.e., not publicly available). Also, whenever a public dataset was utilised, it was generally not maritime-specific. However, recent studies have started a new trend of creating and using publicly available maritime-specific datasets for evaluations, and examples of such public datasets include MARSIM [41], SOLIDS [31], and RadarPWN [59], which have been utilised in [41, 31, 33, 30].
@@
Table IV also reveals several other aspects of method evaluations – not all studies have published widely acknowledged performance metrics like precision, recall, and F1-score which provide a realistic picture of the method performance (for such studies, the “Performance Evaluation” column in Table IV contains no). Also, many studies have not released the implementation of the method (see the “Is Code Publicly Available” column in Table IV) and have not assessed the computational cost of the method (see the “Computational Cost Assessment” column in Table IV).
@@
The following sections provide a more detailed discussion of all papers from Table III. For structuring the discussion, we have closely followed the aforementioned clustering of the selected research papers into four clusters based on the taxonomy presented in Figure 2. The papers on security monitoring architectures and testbeds (with the “Paper Type” column set to A or T in Table III) are covered in Sections V-E and V-F, respectively (note that such papers correspond to the orange cluster in Figure 2). The discussion of the remaining papers on security monitoring methods (with the “Paper Type” column set to M in Table III) has been divided into three subsections by the targeted environment (represented by red, blue, and green clusters in Figure 2) – the papers concerning security monitoring methods for distributed maritime systems and networks are covered in Section V-B, papers on security monitoring methods for navigational systems are covered in Section V-C, and papers on security monitoring methods for maritime CPSs are covered in Section V-D.